# MuLBA: Multidimensional Adaptive Language-based Assessments for Overlapping Mental Health Conditions

### v2

@@ 2. Question Information
We find that two significant factors emerged based on parallel analysis (Horn 1965). The first factor is characterized by strong loadings from measures of depression, anxiety, PTSD, ADHD, and autism, suggesting that it reflects a broad internalizing or emotional distress dimension. The second factor is dominated by high loadings for drug and alcohol use, pointing toward a substance use or externalizing dimension. Conditions such as bipolar disorder, PTSD, and alcohol use disorders exhibit notable loadings on both factors (within 20% of each other), indicating that they share features with both internalizing and externalizing constructs.
@@
We use the mirtCAT [^2] https://CRAN.R-project.org/package=mirtCATLink to mirtCAT., a computerized adaptive testing framework based on mirt to implement adaptive testing. The adaptive testing is done by choosing the most informative question at each turn, determined using D-optimality, a heuristic determined to be optimal for multidimensional IRT in (Chalmers 2016). It maximises the determinant of the Fisher information matrix at each turn of the questions, thus maximizing information gain across all the underlying dimensions.
@@
We compare random ordering of the questions to that of adaptive ordering using MAQuA, and report the Pearson correlations of the estimated scores at each step against the validated clinical scales in Table 4. Further, we also calculate the stability of standard deviation of the estimates at each turn of question-asking, setting the threshold at 0.01 to determine the point at which the estimated mental health scores from MAQuA for each dimension mostly stabilizes and approaches convergence. This metric can also help determine early stopping when deploying MAQuA (Figure 3).
Results
@@
Table 4 shows the when the estimated scores from MAQuA are significantly better than their random counterpart. After the first question, MAQuA seems to consistently jump to improve estimates across all dimensions, whereas in the case of random, the jumps are inconsistent. Based on the correlation curves presented in Figure 3, both random ordering and MAQuA lead to a plateau in score estimation across most of the dimensions, suggesting that modeling multiple conditions with two underlying factors can lead to the model learning the shared semantics across the conditions, regardless of the question-asking strategy used. MAQuA doesn’t always converge better: in the case of ADHD, Autism and OCD, random question-asking is just as good as MAQuA in terms of when convergence (small change in estimated scores in subsewuent turns) or maximum performance is reached. Despite that, MAQuA outperforms random ordered question-asking, especially for conditions like bipolar, alcohol and drug use, depression, anxiety and eating disorder, especially in the first few turns (2-12). This suggests that both the factors, internalizing and externalizing, are being prioritized almost equally when optimizing for information gain over multiple turns of question-asking, which indicates that the chosen model is quite effective in adaptive question-asking while optimizing across ten distinct conditions.
Figure 3: Pearson correlations of MAQuA-estimated scores over the number of questions asked along with their rolling standard deviation of the correlations. The vertical line shows the stability of the estimation based on a threshold for the standard deviation. Our adaptive method consistently stabilizes in at most 50% the number of questions as random.
@@
To determine early stopping criteria and when a question-asking session could be potentially shortened, we also report the stabilization points in Figure 3 by marking the question numbers. Stabilization does not necessarily mean convergence, it might indicate slow ascent as well. The vertical lines indicate the point (nth question) where the rolling standard deviation drops below a threshold. As reported in Table  4, employing the early stopping rule could lead to 508550-85% reduction in the number of questions across all the mental health conditions being evaluated simultaneously. This offers significant potential to save time for both LLM-patient and clinician-patient interactions while reducing overall burden.
Condition 1 8 16 24 32 48
Ques. to
stabilize
Depression
Random .327 .642 .686 .759 .753 .778 42
[…]
Anxiety
Random .332 .612 .647 .690 .694 .730 34
[…]
ADHD
Random .282 .490 .497 .568 .569 .579 41
[…]
Autism
Random .184 .320 .298 .361 .354 .360 43
[…]
Bipolar
Random .202 .420 .390 .452 .454 .473 >> 48
[…]
Eating Disorder
Random .096 .197 .199 .200 .223 .263 >> 48
[…]
OCD
Random .143 .253 .261 .278 .303 .318 39
[…]
Alcohol Use
Random .115 .191 .249 .269 .285 .309 40
[…]
PTSD
Random .236 .474 .468 .502 .499 .526 35
[…]
Drug Use
Random .078 .287 .309 .343 .430 .451 >> 48
[…]
Table 4: Random and MAQuA Scores by condition across the 48 questions, each averaged across 20 runs. The last column shows the number of questions it takes for stabilization.
Conclusion
@@
These results highlight the potential in LLM mental health agents for combining advanced language understanding with psychometrically grounded adaptive testing to overcome limitations of prior approaches that face challenges related to the inconsistency of LLM responses, and are not optimized to reduce patient burden. MAQuA’s effectiveness in modeling transdiagnostic symptom profiles marks an important step toward scalable, interactive, and clinically valid mental health screening tools. Future research should extend and test this framework on real-time conversational settings, improving generalizability in diverse clinical populations.
Appendix A Limitations
This work has several important limitations. First, all participants provided responses in English and were primarily recruited from the UK, Sweden, and the US, which may restrict the applicability of our findings to other languages and cultural settings. Additionally, ADHD is underrepresented in the dataset, limiting the reliability of conclusions related to this condition. Our experiments focus on a fixed set of questions, which helps control the assessment but may miss the full potential of open-ended question selection; exploring more flexible, open-ended approaches remains an important area for future work. While the multitask capabilities of large language models (LLMs) are critical, they were not explicitly explored in this study. Specifically, the effects of reinforcement learning from human feedback (RLHF) or direct preference optimization (DPO) on question sequencing and preferences are unexplored and warrant further investigation.
@@ Appendix A Limitations
Appendix B Details on Multidimensional IRT
A detailed description of the modeling of MIRT is described in Table A.1. Most of the terms and explanations were derived from  Chalmers (2012).
@@ Appendix B Details on Multidimensional IRT
$\boldsymbol{\theta}=(\theta_{1},\theta_{2},\ldots,\theta_{m}),$
@@
The latent traits $\boldsymbol{\theta}$ are typically assumed to follow a multivariate normal distribution:
$\boldsymbol{\theta}\sim\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma}),$
where $\boldsymbol{\mu}$ is the mean vector (often 0) and $\boldsymbol{\Sigma}$ is the covariance matrix capturing correlations among the
@@
The probability that an individual $i$ with latent trait vector $\boldsymbol{\theta}_{i}$ responds correctly (or supports)
@@
$P(u_{ij}=1\mid\boldsymbol{\theta}_{i},\boldsymbol{a}_{j},d_{j})=\frac{1}{1+\exp\left[\left(\boldsymbol{a}_{j}^{\prime}\boldsymbol{\theta}_{i}+d_{j}\right)\right]},$
where $\boldsymbol{a}_{j}$ is the discrimination vector for item $j$, and $d_{j}$ is the difficulty (threshold) parameter.
@@
the probability of responding in category $k$ given latent traits $\boldsymbol{\theta}$ is modeled as:
$P(Y_{j}=k\mid\boldsymbol{\theta})=P(Y_{j}\geq k\mid\boldsymbol{\theta})-P(Y_{j}\geq k+1\mid\boldsymbol{\theta}),$
where each $P(Y_{j}\geq k\mid\boldsymbol{\theta})$ is computed using a multidimensional logistic function involving the
discrimination vector $\boldsymbol{a}_{j}$, latent trait vector $\boldsymbol{\theta}$, and category threshold parameters $b_{jk}$. This approach
@@
Appendix C Dataset Details
@@
Table A.2: The language response question that the participants responded to, along with the type of response and question code. The relevance of the question codes to specific factors is shown in Table A.3.
Appendix D Multi-outcome Model Prediction Structure
We show the difference between how the “ground truth” or validated clinical scores are related to diagnoses as compared to the predicted scores from the best performing multi-outcome model in Figure A.2. Among ten conditions, only three conditions show significant differences – Bipolar, Autism and Alcohol use. While Bipolar and Autism are better captured by the clinical scores, the predicted scores for Alcohol Use actually outperform the clinically validated scores (AUDIT), showing that alcohol use might be better captured when the other conditions are taken into account as well. Moreover, diagnoses could be preliminary and wrong (Mendel et al. 2011), since it could be a proxy for some other mental health condition. This could lead to spurious correlations that are better disambiguated with signals from a multitask model.
Figure A.2: Cohen’s d against the reported diagnoses for our best multitask model against the validated clinical scores (considered ground truth in the modeling). * indicates one being significantly better correlated to diagnosis than the other.
Appendix E Question-level factors loading
@@
After training the multi-outcome models, we found the multitask input aggregation model to perform the best. This model was then fed question-answer pair representations for all the training set as input, and the model inferred scores for each of the question-answer pair across all the users. These scores were then aggregated at a question-level for applying the factor analysis model derived on user-aggregated scores, to understand how much each question contributes to understanding the two factors found to be significant, which is shown in Table A.3. This was used to define the MIRT model for training and adaptive testing.
Factor Question IDs Condition/Symptom
Table A.3: Question-level mapping to factors that was used to specify the MIRT model.
Appendix F Average Score Stabilization Points
Figure A.3 shows the convergence and stabilization points for averaged correlation across all the ten conditions. We find that on an average, our adaptive methods takes about $12$ questions to converge as well as to reach the stabilization point (threshold=0.01=0.01) whereas random question-asking takes $24$, which is twice as much.
Figure A.3: The stabilization points for the correlations averaged across all the 10 mental health conditions.

### v3

@@ 5.2 Factor Analysis
We find that two significant factors emerged based on parallel analysis horn1965rationale. The first factor is characterized by strong loadings from measures of depression, anxiety, PTSD, ADHD, and autism, suggesting that it reflects a broad internalizing or emotional distress dimension. The second factor is dominated by high loadings for drug and alcohol use, pointing toward a substance use or externalizing dimension. Conditions such as bipolar disorder, PTSD, and alcohol use disorders exhibit notable loadings on both factors (within 20% of each other), indicating that they share features with both internalizing and externalizing constructs.
5.3 Adaptive Testing
@@
We use the mirtCAT [^2] https://CRAN.R-project.org/package=mirtCAT, a computerized adaptive testing framework based on mirt to implement adaptive testing. The adaptive testing is done by choosing the most informative question at each turn, determined using D-optimality, a heuristic determined to be optimal for multidimensional IRT in chalmers2016generating. It maximises the determinant of the Fisher information matrix at each turn of the questions, thus maximizing information gain across all the underlying dimensions.
Discretization
Since IRT models operate on discrete, ordered responses, continuous scores obtained from regression cannot be directly used as inputs. Instead, these continuous predictions must be discretized into $k$ ordinal levels to allow the model to estimate category response functions. This discretization enables the IRT model to capture how each level probabilistically relates to underlying latent dimensions, following standard practices in multidimensional IRT modeling reckase200618. We select the maximum number of discrete levels that ensures sufficient representation across the training data, following the heuristics in prior work on discretizing language-based scores for IRT varadarajan-etal-2024-alba. We discretize scores into four ordinal levels based on the quartile thresholds of the model outputs. Using greater granularity leads to data sparsity, as certain levels would be underrepresented, reducing model stability and interpretability.
We compare random question ordering to adaptive sequencing using MAQuA, and further benchmark these against GPT-4, one of the most popular LLMs previously studied for its mental health estimation capabilities and ability to handle long contexts (moell2024comparing; gpt4openai2023; ganesan2024explaining). At each step, we report the Pearson correlations between estimated scores and validated clinical scales, summarized in Figure 2 and Table D. The prompt can be found in Table A.5.
@@
Further, we also compare the ordering and estimation capabilities of gpt4 to showcase the efficacy of IRT in adaptive question-asking. We also calculate the stability of standard deviation of the estimates at each turn of question-asking, setting the threshold at 0.01 to determine the point at which the estimated mental health scores from MAQuA for each dimension mostly stabilizes and approaches convergence. This metric can also help determine early stopping when deploying MAQuA (Figure 2).
Figure 2: Pearson correlations of MAQuA-estimated scores over the number of questions asked along with their rolling standard deviation of the correlations. The vertical line shows the stability of the estimation based on a threshold for the standard deviation. Our adaptive method consistently stabilizes in at most 50% the number of questions as random. While GPT-4 shows promise in its estimation capabilities of factor 2 (externalizing disorders) - substance use, alcohol use, eating disorder. For more details please see §D .
6 Results
@@
Multi-outcome Model Prediction Structure
We show the difference between how the “ground truth” or validated clinical scores are related to diagnoses as compared to the predicted scores from the best performing multi-outcome model in Figure 3. Among ten conditions, only three conditions show significant differences – Bipolar, Autism and Alcohol use. While Bipolar and Autism are better captured by the clinical scores, the predicted scores for Alcohol Use actually outperform the clinically validated scores (AUDIT), showing that alcohol use might be better captured when the other conditions are taken into account as well. Moreover, diagnoses could be preliminary and wrong mendel2011confirmation, since it could be a proxy for some other mental health condition. This could lead to spurious correlations that are better disambiguated with signals from a multitask model.
Figure 3: Cohen’s d against the reported diagnoses for our best multitask model against the validated clinical scores (considered ground truth in the modeling). * indicates one being significantly better correlated to diagnosis than the other.
@@
Figure 2 shows the when the estimated scores from MAQuA are significantly better than their random counterpart, as well as GPT-4. After the first question, MAQuA seems to consistently jump to improve estimates across all dimensions, whereas in the case of random, the jumps are inconsistent. Based on the correlation curves presented in Figure 2, all the three methods can lead to a plateau in score estimation across most of the dimensions, suggesting that modeling multiple conditions can lead to the model learning the shared semantics across the conditions, regardless of the question-asking strategy used. MAQuA doesn’t always converge better: in the case of ADHD, Autism and OCD, random question-asking is just as good as MAQuA in terms of when convergence (small change in estimated scores in subsewuent turns) or maximum performance is reached. Despite that, MAQuA outperforms random ordered question-asking, especially for conditions like bipolar, alcohol and drug use, depression, anxiety and eating disorder, especially in the first few turns (2-12). This suggests that both the factors, internalizing and externalizing, are being prioritized almost equally when optimizing for information gain over multiple turns of question-asking, which indicates that the chosen model is quite effective in adaptive question-asking while optimizing across ten distinct conditions.
In contrast, GPT-4’s estimation and question-asking behavior differ. Despite being prompted to assess all ten conditions simultaneously (rather than focusing on two factor scores), GPT-4 notably excels over the IRT-based methods in handling externalizing factors: specifically eating disorder, substance use, and alcohol use conditions. This advantage may stem from limitations inherent in the item response theory model, particularly regarding its representation of comorbidity effects within each data subset. Evidence supporting this includes the regression model’s performance, which approximates GPT-4’s correlation levels (Table 1). These findings suggest that a hybrid modeling approach could more effectively capture both internalizing and externalizing aspects.
@@
To determine early stopping criteria and when a question-asking session could be potentially shortened, we also report the stabilization points in Figure 2 by marking the question numbers. Stabilization does not necessarily indicate peak performance, it indicates slower ascent after that point. The vertical lines indicate the point (nth question) where the rolling standard deviation drops below a threshold. As reported in Table  A.3, employing the early stopping rule could lead to 508550-85% reduction in the number of questions across all the mental health conditions being evaluated simultaneously. Figure 2 shows stabilization points for individual and averaged correlation across all the ten conditions. We find that on an average, MAQuA takes about $12$ questions to reach the stabilization point (threshold=0.01=0.01) whereas random question-asking takes $24$ and GPT-4 doesn’t converge at $48$ questions. This offers significant potential to save time for both LLM-patient and clinician-patient interactions while reducing overall burden.
7 Conclusion
@@
These results highlight the potential in LLM mental health agents for combining advanced language understanding with psychometrically grounded adaptive testing to overcome limitations of prior approaches that face challenges related to the inconsistency of LLM responses, and are not optimized to reduce patient burden. MAQuA’s effectiveness in modeling transdiagnostic symptom profiles marks an important step toward scalable, interactive, and clinically valid mental health assessment tools. Future research should extend and test this framework on real-time conversational settings, improving generalizability in diverse clinical populations.
Limitations
This work has several important limitations. First, all participants provided responses in English and were primarily recruited from the UK, Sweden, and the US, which may restrict the applicability of our findings to other languages and cultural settings. Additionally, ADHD is underrepresented in the dataset, limiting the reliability of conclusions related to this condition.
We use a fixed set of questions to maintain control over the assessment and prevent large language models from hallucinating or generating harmful, misleading questions. Although this limits open-ended question selection, it is an important safeguard for participant safety and assessment reliability. Future work will explore flexible approaches with robust hallucination mitigation strategies. While the multitask capabilities of large language models (LLMs) are critical, they were not explicitly explored in this study. Specifically, the effects of reinforcement learning from human feedback (RLHF) or direct preference optimization (DPO) on question sequencing and preferences are unexplored and warrant further investigation.
@@
Ethical Considerations
Participants were compensated hourly for their time, ensuring fair recognition of their contribution and respecting the principle of voluntary and informed participation. All responses were anonymized to protect participant confidentiality, and the data were stored securely on protected servers to safeguard privacy and comply with data protection standards. The study received approval from the Institutional Review Board (IRB) [redacted], assuring that appropriate ethical oversight was in place throughout the research process.
MAQuA is designed to reduce the burden of mental health assessments and address diagnostic inconsistencies, thereby promoting ethical principles of minimizing participant fatigue and improving assessment accuracy. By combining multi-outcome modeling with multidimensional item response theory (MIRT), MAQuA selects the most informative questions adaptively, reducing assessment length by up to 85% while maintaining psychometric validity, which supports participant well-being and respects their time.
MAQuA represents a methodological integration of advanced natural language processing and psychometric measurement theory, grounded in ethical commitments to scientific integrity, transparency, and reproducibility. The release of the accompanying questionnaire-driven dataset encourages further validation and responsible innovation, ensuring that AI tools in mental health are developed with accountability and sustained patient trust.
Appendix A Details on Multidimensional IRT
A detailed description of the modeling of MIRT is described in Table A.1. Most of the terms and explanations were derived from  chalmers2012mirt.
@@ Appendix A Details on Multidimensional IRT
$\bm{\theta}=(\theta_{1},\theta_{2},\ldots,\theta_{m}),$
@@
The latent traits $\bm{\theta}$ are typically assumed to follow a multivariate normal distribution:
$\bm{\theta}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma}),$
where $\bm{\mu}$ is the mean vector (often 0) and $\bm{\Sigma}$ is the covariance matrix capturing correlations among the
@@
The probability that an individual $i$ with latent trait vector $\bm{\theta}_{i}$ responds correctly (or supports)
@@
$P(u_{ij}=1\mid\bm{\theta}_{i},\bm{a}_{j},d_{j})=\frac{1}{1+\exp\left[\left(\bm{a}_{j}^{\prime}\bm{\theta}_{i}+d_{j}\right)\right]},$
where $\bm{a}_{j}$ is the discrimination vector for item $j$, and $d_{j}$ is the difficulty (threshold) parameter.
@@
the probability of responding in category $k$ given latent traits $\bm{\theta}$ is modeled as:
$P(Y_{j}=k\mid\bm{\theta})=P(Y_{j}\geq k\mid\bm{\theta})-P(Y_{j}\geq k+1\mid\bm{\theta}),$
where each $P(Y_{j}\geq k\mid\bm{\theta})$ is computed using a multidimensional logistic function involving the
discrimination vector $\bm{a}_{j}$, latent trait vector $\bm{\theta}$, and category threshold parameters $b_{jk}$. This approach
@@
Appendix B Dataset Details
@@
Table A.2: The language response question that the participants responded to, along with the type of response and question code. The relevance of the question codes to specific factors is shown in Table 5.
Appendix C Question-level factors loading
@@
After training the multi-outcome models, we found the multitask input aggregation model to perform the best. This model was then fed question-answer pair representations for all the training set as input, and the model inferred scores for each of the question-answer pair across all the users. These scores were then aggregated at a question-level for applying the factor analysis model derived on user-aggregated scores, to understand how much each question contributes to understanding the two factors found to be significant, which is shown in Table 5, and visually shown as a wordcloud in Figure A.2. This was used to define the MIRT model for training and adaptive testing.
Figure A.2: Question texts loading on to the two factors.
Appendix D Stabilization Points
The stabilization points are given in a table along with the improvement over the random ordering and GPT-4 in Table A.3.
Condition 1 8 16 24 32 48
Ques. to
stabilize
Depression
GPT-4 -.001 .375 .518 .618 .637 .608 >> 48
Random .327 .642 .686 .759 .753 .778 42
[…]
Anxiety
GPT-4 -.029 .293 .461 .527 .565 .587 >> 48
Random .332 .612 .647 .690 .694 .730 34
[…]
ADHD
GPT-4 - .248 .419 .447 .513 .486 >> 48
Random .282 .490 .497 .568 .569 .579 41
[…]
Autism
GPT-4 - .180 .324 .441 .496 .502 >> 48
Random .184 .320 .298 .361 .354 .360 43
[…]
Bipolar
GPT-4 - .195 .323 .398 .474 .448 >> 48
Random .202 .420 .390 .452 .454 .473 >> 48
[…]
Eating Disorder
GPT-4 - .213 .358 .451 .479 .468 >> 48
Random .096 .197 .199 .200 .223 .263 >> 48
[…]
OCD
GPT-4 - .163 .246 316 .307 .336 >> 48
Random .143 .253 .261 .278 .303 .318 39
[…]
Alcohol Use
GPT-4 - .343 .425 .525 .535 .537 >> 48
Random .115 .191 .249 .269 .285 .309 40
[…]
PTSD
GPT-4 - .220 .330 .407 .438 .424 >> 48
Random .236 .474 .468 .502 .499 .526 35
[…]
Drug Use
GPT-4 - .240 .489 .530 .516 .521 >> 48
Random .078 .287 .309 .343 .430 .451 >> 48
[…]
Table A.3: Random and MAQuA Scores by condition across the 48 questions, each averaged across 20 runs. The last column shows the number of questions it takes for stabilization.
Carefully read the entire transcript, considering both explicit answers and patterns, emotional tone, and recurring themes throughout the conversation.
Rely exclusively on the transcript; do not infer information not present or speculate beyond what is written.
For each mental health dimension below, provide only the mental health scores requested, nothing more:
A single numerical score, adhering to the questionnaire’s scale (see below for ranges).
If insufficient information is present to make an informed rating for a dimension, give a score of 0.
Mental health dimensions and corresponding standardized rating scales:
Depression score: PHQ-9 (range: 0–27)
Anxiety score: GAD-7 (range: 0–21)
Mood_Disorder score: MDQ (range: 0–15)
Obsessive_Compulsive Disorder score: Y-BOCS (range: 0–40)
Autism score: RAADS (range: 0–42)
Eating_Disorder score: EDE-QS (range: 0–72)
PTSD score: NSESSS (range: 0–36)
Alcohol_Use score: AUDIT (range: 0–40)
Substance_Use score: DUDIT (range: 0–44)
ADHD score: ASRS Part A (range: 0–24)
Table A.4: Prompt for score estimation with Llama3.2 1B model.
You are an expert conversational agent conducting a mental health assessment.
Your goal is to select, one at a time, the most informative question from a provided list to accurately score ten standardized mental health dimensions.
Use prior conversation context to guide your choices. After each patient response, reassess and select the next optimal question.
If no further questions will improve scoring or none remain, output final scores for each dimension.
Never answer the questions yourself—only select a question to ask or output the final scores.
<interview_history>
This section contains the interview history or conversational context so far.
Use this information to inform your next question selection.
If this is empty, you are at the start of the interview.
</interview_history>
<candidate_questions>
This section contains the list of questions left to be asked at the current turn.
Each question is labeled with a unique code and text.
Select the most informative question from this list to ask next, unless you determine that the assessment should stop.
</candidate_questions>
Output format is as below. Only output in this JSON format with the mentioned keys, don’t output anything else:
question_code: a single question code from candidate_questions to ask next, or -1 if stopping,
depression_score: X/100,
anxiety_score: X/100,
mood_disorder_score: X/100,
ocd_score: X/100,
autism_score: X/100,
eating_disorder_score: X/100,
ptsd_score: X/100,
alcohol_use_score: X/100,
substance_use_score: X/100,
adhd_score:X/100
Table A.5: Adaptive Question-asking prompt for GPT-4